# Skyrise: Exploiting Serverless Cloud Infrastructure for Elastic Data Processing

### v1

@@ 1 Introduction
Previous work shows the viability of systems built on serverless infrastructure for infrequent and interactive analytical workloads [22, 23]. Ana-logously, this workload category is also targeted by serverless databases that integrate automatic and elastic scalability at the system level [24, 25, 26].
@@
    We propose adaptive techniques for robust query performance on performance-variable cloud resources. Specifically, we describe the pre-warming of serverless resources and the re-triggering of straggling serverless query workers.
@@ 2.1 Function as a Service Platforms
Serverless function platforms like AWS Lambda are commonly realized by combining client-facing load balancing, cluster management, and functions / workers for which the system communicates through eventing. The cloud function user submits application binaries to be executed as ZIP archives or container images through a load balancing frontend service, coordinating function invocations. The request frontend checks the function metadata with the admission (control) service of the cluster management and—if the request does not exceed the user’s quota of concurrent function executions—it issues sandboxed environments from a pool of workers for the execution of the binary. Sandboxes are virtualized resources that are reused to place incoming requests, making them small, short, and ephemeral tasks.
Current vendors of FaaS platforms limit their configurations as shown in Table 1 with function sizing based on memory capacity that determines the number of virtual CPUs. Notably, function sizes are an order of magnitude smaller than VMs such as EC2 instances. For serverless functions hour-long lifetimes, persistent state, and direct communication are prohibited, making the functions small, short, ephemeral tasks that are easy to manage.
@@ 2.2 Serverless Storage Services
From the cost perspective, S3 Standard is the most economical storage choice with the lowest static storage costs, no transfer costs, but the highest costs for number of requests. In contrast, S3 Express halves the request costs, but introduces transfer costs and seven times higher storage costs. The other storage options have factor around 10 higher storage costs compared to S3 Standard and either demand costs per request for DynamoDB or data transfers in case of EFS.
@@ 3 System Architecture
In this section, we introduce the Skyrise serverless query processor. We provide an overview of its architecture and describe the core components, which fall into three layers: Query compilation, query execution, and storage management.
@@ 3.2 Query Compilation
The query compilation pipeline in Skyrise takes a SQL query and generates a physical plan for distributed data-parallel execution by serverless query workers. The pipeline consists of multiple steps and is depicted in Figure 2. The first set of steps resemble the SQL frontend (in green). This component parses a query string and builds an abstract syntax tree (AST) that is further transformed into a logical plan for this query (LQP). Thereby, semantic types of columns in referenced tables are validated against an external database catalog. Then, a rule-based optimizer framework applies a range of conventional heuristics-based rules to the plan structure. These rules include predicate pushdown, subquery flattening, and join ordering. They only need simple statistics and are oblivious of Skyrise’s distributed execution environment. Next, the query optimizer (in blue) employs another set of rules to produce a physical query plan (PQP). The rules map logical operators to physical operators, e.g., a repartition join vs. a broadcast join. They identify query pipeline breakers in the plan to introduce shuffle points as necessary. And finally, they prepare the plan for serverless execution. This involves deciding on the number and size of workers based on the input data partitions, as well as picking a strategy for shuffling, e.g., staging to reduce cost [22, 23] or tiering to reduce runtime [38]. The final artifacts of query compilation are the per-pipeline PQPs parameterized with the input and output data partitions for the individual workers. These plan fragments are then scheduled for execution.
@@ 3.3 Query Execution
The execution of a query in Skyrise is orchestrated by the query coordinator (cf. Figure 1). Once a query is compiled into a set of interdependent pipeline PQPs, the coordinator schedules these PQPs stage-wise based on their dependencies. For each fragment of a pipeline, the coordinator invokes a serverless function running an instance of the Skyrise query worker. When the number of fragments is large and function invocation can dominate query runtime, the coordinator employs a two-level invocation procedure [22]. For $W$ fragments, it invokes $\sqrt{W}$ workers, each with a list of $\sqrt{W}$ fragments that in turn invoke $\sqrt{W}-1$ workers before executing their own fragment. The workers and the coordinator both invoke cloud functions asynchronously, decoupling their lifetime from the lifetime of their direct child functions. A function request is a query fragment serialized in JSON. Upon startup, a query worker first deserializes its invocation payload into the PQP form. Then, it schedules and executes the operators accordingly. The worker fetches its input data partitions from shared storage (cf. Section 3.4) and transforms them into an in-memory columnar format. From there, the execution follows a push-based, vectorized model. The query operators push their results in batches to downstream operators. Operators with no interdependencies run in parallel. Operators further exploit the thread-level parallelism available in serverless function workers. The final operator in a fragment materializes the results to shared storage. Before the worker quits, it sends a response message to a queue that is polled by the coordinator. The response contains the result location and execution statistics that can be used for adaptive execution behavior and query billing. When the coordinator has received responses from all workers in a stage, it proceeds with scheduling the next stage until the query finishes. Upon query completion, the coordinator responds to the user with the location of the final query result.
@@ 3.4 Storage Management
Figure 3: Storage I/O stack of the Skyrise query worker including the interactions between query operators and cloud storage.
@@
The input handler splits large object storage requests from the execution layer into several smaller requests, which are processed in parallel. The byte offsets of the small requests are chosen based on the object’s PAX layout. This way only relevant columns and rows are fetched. The result of each request is stored in a dedicated buffer. Straggling requests are retriggered after a timeout based on their size [23]. Once buffered, the results are deserialized and decompressed batch-wise and on demand, keeping the memory footprint of the I/O stack small. The output handler takes data batches from the execution layer and serializes, compresses, buffers them as they arrive. Once the query worker’s result is complete, it is materialized to remote storage as a single object.
@@ 4 Experimental Evaluation
In this section, we conduct a series of experiments to answer the following questions about the FaaS-based serverless query processor Skyrise:
  • How does Skyrise’s query performance compare to other cloud-based data processing systems?
@@ 4.1 Setup and Methodology
We conduct all experiments on AWS in the region us-east-1 [42] and in the time frame of August to October 2024. For our experiments on Skyrise, we deploy ARM-based Lambda functions. Our driver has low resource requirements and runs continuously on an c6g.xlarge instance across experiments. All other compute resources are newly created for each experiment configuration and repetition. As cloud data processing systems, we consider AWS Athena [24], Snowflake [43], and Lambada [22]. We use queries and datasets from the TPC-H [44] benchmark. The tables are partitioned into 180 MB Parquet files using ZSTD compression and are stored on S3. We employ the standard generators and do not partition or sort on any specific keys. The lineitem table is 177.4 GiB in total and the orders table 44.9 GiB.
@@ 4.2 End-to-End Workloads
To assess the performance and cost efficiency of Skyrise, we compare it against several state-of-the-art approaches: The academic research prototype Lambada [22] and the commercial offerings Snowflake and Athena. Lambada (see Section 5) is a research prototype for FaaS-based query processing. We report the numbers published in [22]. Snowflake is a commercial cloud-based data warehouse offering. Typically, users allocate several warehouse instances. Recently, Snowflake added serverless tasks which are mainly aimed at data maintenance. Serverless tasks are executed on warehouses provisioned by Snowflake and are billed per second. Serverless tasks are created with an initial warehouse size provided by the user. After a few executions, Snowflake automatically determines the warehouse size for each tasks. For stable measurements, we created new tasks per query execution and set the initial size to 2xlarge (the configuration with the best price-performance ratio). We loaded the TPC-H dataset as external tables. We use Snowflake version 8.38.3 and take costs provided by Snowflake’s monitoring tools (with a standard configuration costing USD 2 per compute credit). Athena is a serverless query processing offering by Amazon. Similar to the Snowflake setup, we created external tables based on the TPC-H dataset. For query execution, Athena uses Trino (previously Presto). Resources used for queries are automatically determined by Athena and cannot be adapted or observed by users. The determine costs per query, we use the reported bytes scanned by Athena and the current pricing model of USD 5 per terabyte of data scanned.
We evaluate the three scan-heavy TPC-H queries 1, 6, and 12 (scale factor 1000), which are well-suited for serverless execution as they are not dominated by shuffling operators. We measure runtimes and costs per query execution, reporting the median of ten executions.
@@ 4.2.1 Query Latency
Users with ad-hoc analytical query workloads want interactive query performance. We present our results in Figure 4. We see a similar pattern for all queries. The fastest system is in all cases Lambada, followed by Skyrise, which is on average 30 % slower. The two commercial systems Athena and Snowflake are slower for evaluated queries with Athena being 5×5$\times$5 × faster than Snowflake in average.
We assume the reason for the performance difference between Lambada and Skyrise to be the dataset that was used for the Lambada experiments. This dataset differs in two significant ways. First, the string columns in the dataset were replaced with integers. Second, the lineitem relation is sorted on the l_shipdate attribute, which enables partition pruning for the employed queries.
We observe in separate experiments that both commercial systems perform better on smaller scale factors in their serverless modes, but their performance degrades on terabyte-scale queries.
@@ 4.2.2 Monetary Cost
Figure 5 shows the average costs per query. The pattern is comparable to the latency results. Lambada has the lowest costs per query, followed by Skyrise. Athena and Snowflake follow with Athena being 7×7$\times$7 × cheaper in average.
To conclude, Skyrise achieves comparable latencies and costs as the state-of-the-art system Lambada. The commercial system Athena closely follows the research systems, which is noteworthy given that the non-commercial systems Lambada and Skyrise only incur infrastructure costs. In our tested scenario, the recent Snowflake serverless tasks offering cannot compete with Athena.
@@ 4.3 Elasticity
To evaluate the elasticity of Skyrise, we run TPC-H queries 1 and 6 for five different scale factors, ranging from 1 to 10 000. We run the queries cold with 15 minute intervals between runs to avoid any caching or provisioning effects in the AWS infrastructure. The results are shown in Figure 6. We report the aggregated runtimes of both queries. We observe that larger scale factors, especially 1 000 and 10 000 are generally more prone to stragglers in function invocations and storage requests. This is particularly the case for the 10 000 configuration, which runs with a maxmimum of 2,500 query workers and suffers from IOPS limits in S3. Generally, without any upfront provisioning or query load, Skyrise automatically scales the resources with the size of the query.
@@ 5.1 FaaS-based Data Processing
After its first public availability in 2014, FaaS technology was first prominently picked up for cloud computing by Jonas et al. who developed PyWren [16]. PyWren is a prototypical implementation of a programming model for disaggregated cloud computing environments using stateless functions.
While PyWren more generally targeted cloud applications, the potential of cloud functions for data analytics was shown by Kim et al. with Flint [45] and later Pu et al. with Locus [46]. Flint is a rewrite of Apache Spark’s execution layer using serverless building blocks like workers and disaggregated storage. In addition, Locus proposed a shuffling operator and intermediate data exchange using AWS ElasticCache and S3.
For data processing on serverless infrastructure, Locus was extended by Zhang et al. with a serverless SQL engine [47], called Caerus, focusing on analytical query processing atop AWS Lambda and cost-efficiency. The Lambada system [22] spawns static cluster for all query pipelines.
The shuffling approaches from PyWren, Flint, and [48] were further enhanced by Lambada and Starling [23] through stage shuffles, reducing shuffle I/Os via ephemeral storage. In addition Starling addresses robust performance processing data on S3 through pipelining and solutions for data skew and stragglers. Tiered shuffle was designed in Locus, Cackle [49]/Pixels [50], Caerus /Jiffy [51] and Pocket [52]. Boxer [53] shuffles via NAT and Flock [54] for streaming utilizes function invocation payloads [54].
@@
Combining stateful and stateless compute is an interesting new area that Pixels [50] and Cackle [49] introduced as hybrid compute.

### v2

@@ 1 Introduction
Previous work shows the viability of systems built on serverless infrastructure for infrequent and interactive analytical workloads [22, 23]. These vary over time and cannot be accurately predicted. One example is a few-shot, resource-intensive query on cold data. This workload category is also targeted by serverless databases that integrate elasticity at the system level [24, 25, 26].
@@
    We propose adaptive techniques for robust query performance on performance-variable cloud resources. Specifically, we describe the re-triggering of straggling query workers.
@@ 2.1 Function as a Service Platforms
Function as a service (FaaS) platforms, such as AWS Lambda are realized by combining client-facing load balancing, cluster management, and functions for which the system communicates through eventing. The cloud function user submits application binaries to be executed as ZIP archives or container images through a frontend service, coordinating function invocations. The request frontend checks the function metadata with the admission (control) service of the cluster management, and—if the request does not exceed the user’s quota of concurrent function executions—it issues sandboxed environments from a pool of hosts for the execution of the binary.
@@
Current FaaS platforms limit function configuration as shown in Table 1. Sizing is based on memory capacity, which determines the number of virtual CPUs [29]. Notably, function sizes are an order of magnitude smaller than VMs, such as EC2 instances [30]. Functions are short-lived and cannot maintain state beyond their lifetime. They can only communicate indirectly through cloud storage. In addition, they are prone to stragglers, e.g., during initialization or when interacting with remote storage.
@@ 2.2 Serverless Storage Services
From the cost perspective, S3 Standard is the most economical storage choice with the lowest static storage costs, no transfer costs, but the highest costs for the number of requests. In contrast, S3 Express halves the request costs, but introduces transfer costs and seven times higher storage costs. The other storage options have around 10$\times$ higher storage costs compared to S3 Standard and either demand costs per request for DynamoDB or data transfers in case of EFS.
@@ 3 System Architecture
In this section, we introduce the Skyrise serverless query processor [36]. We provide an overview of its architecture and describe the core components, which fall into three layers: Query compilation, query execution, and storage management.
@@ 3.2 Query Compilation
Figure 3: Query execution plan for TPC-H Q12.
The query compilation pipeline in Skyrise takes a SQL query and generates a physical plan for distributed data-parallel execution by serverless query workers. The pipeline consists of multiple steps and is depicted in Figure 2. The first set of steps resembles the SQL frontend (in green). This component parses a query string and builds an abstract syntax tree (AST) that is further transformed into a logical plan for this query (LQP). Thereby, semantic types of columns in referenced tables are validated against an external database catalog. Then, a rule-based optimizer framework applies a range of conventional heuristics-based rules to the plan structure [41]. These rules include predicate pushdown, subquery flattening, and join ordering. They only need simple statistics and are oblivious of the distributed serverless execution environment of Skyrise. Next, a physical optimizer (in blue) employs another set of rules to produce a physical query plan (PQP). The rules map logical operators to physical operators, e.g., a join to a repartition vs. broadcast join. They identify query pipeline breakers in the plan to introduce shuffle points as necessary. And finally, they prepare the plan for serverless execution. The number of workers per pipeline is based on the total input size and the network burst capacity per function, as determined in prior work [42]. The strategy for shuffling takes the request rate limits of object storage into account and reduces runtime via tiering to hotter storage [43]. Since function workers are stateless, the optimizer does not consider cache affinity, and thus data locality. Moreover, the optimizer does not distinguish cold vs. warm starts of functions, because cold starts are negligible (cf. Table 2) and only occur in the initial query stage. Subsequent stages reuse the cached function binaries. The final artifacts of query compilation are the per-pipeline PQPs parameterized with the input and output partitions for the individual workers. An example query plan with its pipelines is shown in Figure 3.
@@ 3.3 Query Execution
The execution of a query in Skyrise is orchestrated by the query coordinator (cf. Figure 1). Once a query is compiled into a set of interdependent pipeline PQPs, the coordinator schedules these PQPs stage-wise based on their dependencies. For each fragment of a pipeline, the coordinator invokes a serverless function running an instance of the Skyrise query worker. When the number of fragments is large and function invocation can dominate query runtime, the coordinator employs a two-level invocation procedure [22] to parallelize the cluster startup. For $W$ fragments, it invokes $\sqrt{W}$ workers, each with a list of $\sqrt{W}$ fragments that in turn invoke $\sqrt{W}-1$ workers before executing their own fragment. The workers and the coordinator both invoke cloud functions asynchronously, decoupling their lifetime from the lifetime of their direct child functions. Functions are neither invoked speculatively before query arrival nor are they kept running between stages to avoid the cost of idling resources. A function request is a query fragment serialized in JSON. Upon startup, a query worker first deserializes its invocation payload into the PQP form. Then, it schedules and executes the operators accordingly. The worker fetches its input data partitions from shared storage (cf. Section 3.4) and transforms them into an in-memory columnar format. Since worker functions cannot persist data locally, this process is repeated for recurring queries. From there, the execution follows a push-based, vectorized model. The query operators push their results in batches to downstream operators. Operators with no interdependencies run in parallel. Operators further exploit the thread-level parallelism available in serverless function workers. The final operator in a fragment materializes the results to shared storage. Before the worker quits, it sends a response message to a queue that is polled by the coordinator. The response contains the result location and execution statistics that can be used for adaptive execution behavior and query billing. When the coordinator has received responses from all workers in a stage, it proceeds with scheduling the next stage until the query finishes. Upon query completion, the coordinator responds to the user with the location of the final query result. Skyrise offers robust query performance on its fine-grained resources, i.e., many small tasks and requests, that may fail individually. To accomplish this, Skyrise employs a variety of adaptive techniques. Skyrise’s coordinator tracks the progress of workers, which can encounter failures, e.g., because of code issues, data skew, or transient cloud infrastructure errors. Based on query context and runtime statistics, the coordinator differentiates between the failure categories. Then, the coordinator restarts individual workers, reassigns fragments to more workers, or aborts the query. Retriggering workers mid-query is practically viable, because it takes milliseconds. It does not invalidate (intermediate) query results, because the worker function is idempotent and only ever writes a single deterministic output file to shared storage. Racing worker functions merely overwrite their (identical) results. Aborted queries can be continued from any complete stage result, since they resemble checkpoints. Query failures can be also investigated post-mortem [44].
@@ 3.4 Storage Management
Figure 4: Storage stack of the Skyrise query worker including the interactions between query operators and cloud storage.
@@
The input handler splits large object storage requests from the execution layer into several smaller requests, which are processed in parallel. The byte offsets of the small requests are chosen based on the object’s PAX layout. In this way, only relevant columns and rows are fetched. The result of each request is stored in a dedicated buffer. Straggling requests are retriggered aggressively after a short timeout. Once buffered, the results are deserialized and decompressed batchwise and on demand, keeping the memory footprint of the I/O stack small. The output handler takes data batches from the execution layer and serializes, compresses, and buffers them as they arrive. Once the query worker’s result is complete, it is written to shared cloud storage as a single object.
All objects resulting from a query pipeline combined represent a (intermediate) query result. Query results are referenced in a central registry with an identifier that is generated during query compilation. In particular, the identifier is a hash on the PQP after logical optimization, yet before most of the physical optimization. This allows matching semantically equivalent results independent of physical execution properties, such as the number and size of the workers producing the result. Before the coordinator schedules a planned query pipeline, it checks the result registry for a respective entry. If the entry exists, the coordinator skips the pipeline and repeats this process for the depending query pipelines. If there is no entry, the query pipeline is scheduled for execution, and its result is registered when complete. This cache makes Skyrise more performant and cost-efficient for workloads with recurring queries.
@@ 4 Experimental Evaluation
In this section, we conduct a series of experiments to answer the following questions about our FaaS-based Skyrise query processor:
  • How does Skyrise’s query performance compare to other cloud-based data analysis systems?
@@ 4.1 Setup and Methodology
We conduct all experiments on AWS in the region us-east-1 [48] in the time frame of August 2024 to January 2025. For our experiments on Skyrise, we deploy ARM-based Lambda functions [49] and deactivate the query result cache. Our driver has low resource requirements and runs on an c6g.xlarge instance [30] across experiments. All other compute resources are newly created for each experiment configuration and repetition. As cloud data analysis systems, we consider AWS Athena [24], Snowflake [50], and Lambada [22]. We use queries and datasets from the TPC-H [51] benchmark. The tables are partitioned into Parquet files using ZSTD compression and are stored on S3. We employ the standard generators and do not partition or sort on any specific keys. For TPC-H scale factor 1,000, the lineitem table totals 177.4 GiB and the orders table 44.9 GiB.
@@ 4.2 End-to-End Workloads
To assess the performance and cost efficiency of Skyrise, we compare it against several state-of-the-art approaches: The academic research prototype Lambada [22] and the commercial offerings Snowflake and Athena. Lambada (see Section 5) is a research prototype for FaaS-based query processing. We report the numbers published in [22]. Snowflake is a commercial cloud-based data warehouse offering. Typically, users allocate several warehouse instances. Recently, Snowflake added serverless tasks which are mainly aimed at data maintenance. Serverless tasks are executed on warehouses provisioned by Snowflake and are billed per second. Serverless tasks are created with an initial warehouse size provided by the user. After a few executions, Snowflake automatically determines the best warehouse size for each tasks. For stable measurements, we created new tasks for each query execution and set the warehouse size to 2xlarge (the configuration with the best price-performance ratio). We evaluated five setups for Snowflake (SF): the default of loading data into the proprietary format (Server/Intern), as external tables from warehouses (Server/Extern) and serverless tasks (Serverless/Extern), and as Iceberg tables from warehouses (Server/Iceberg) and serverless tasks (Serverless/Iceberg), respectively. Both the external tables and the Iceberg tables link the S3-resident Parquets files and execute directly on them. The main difference between both setups is that Snowflake stores meta data for Iceberg tables. As Athena, Lambada, and Skyrise do not store any metadata on the table data, we consider the external table setup to be the most relevant for a fair comparison. We use Snowflake version 8.38.3 and take costs provided by Snowflake’s monitoring tools (with a standard configuration costing \$2 per compute credit). For the warehouse-based Snowflake variants, we multiply the credit price with the query runtime. Athena is a serverless query processing offering by Amazon. Athena builds on Trino (previously Presto [52]). Similar to the setup for Snowflake, we created external tables based on the TPC-H dataset. Resources used for queries are automatically determined by Athena and cannot be adapted or observed by users. For the determined costs per query, we use the reported bytes scanned by Athena and the current pricing model of \$5 per terabyte of data scanned.
We run the three scan-heavy TPC-H queries 1, 6, and 12 (scale factor 1,000), which are well-suited for serverless execution as they are not dominated by shuffling operators. We measure runtimes and costs per query execution, reporting the median of five executions.
@@ 4.2.1 Query Latency
Users with ad-hoc analytical query workloads want interactive response times. We present our results in Figure 5. The overall fastest system is standard Snowflake with its internal data format. It is the only system that does not need to read data from S3. From the remaining systems, the fastest for Q1 and Q6 is Lambada, followed by Skyrise, which is on average 30 % slower. Both systems share key architecture and design features. They exploit the scalability of their underlying FaaS platform and efficiently handle columnar file formats on remote cloud storage. For Q12, the two Snowflake variants working with Iceberg tables are faster than the academic prototypes. They are followed by Athena, which is 23×2-3$\times$2 - 3 × slower in our evaluation. Comparing the external table setups of Snowflake, Athena is 58×5-8$\times$5 - 8 × faster on average. The QaaS systems scale out their resources conservatively in serverless mode, which degrades performance for infrequent terabyte-scale queries. Specifically for Snowflake, the external table setups appear to miss a range of statistics-based optimizations as compared to Iceberg tables.
We hypothesize that the reason for the performance difference between Lambada and Skyrise is the dataset for the Lambada experiments [22]. This dataset differs in two significant ways from ours. First, the string columns are replaced with integers, reducing the time it takes to load and parse them. Second, the lineitem relation is sorted on the l_shipdate attribute, which enables partition pruning for the evaluated queries.
@@ 4.2.2 Monetary Cost
Figure 6 shows the average costs per query. The pattern is comparable to the latency results. Lambada has the lowest costs per query, followed by Skyrise and the Snowflake variants with internal and Iceberg tables. Athena and Snowflake with external tables follow with Athena being 7×7$\times$7 × cheaper on average.
Skyrise achieves comparable latencies and costs as the state-of-the-art system Lambada. The commercial systems Athena and Snowflake (again with internal and Iceberg tables) closely follow the research systems, which is noteworthy given that the non-commercial systems Lambada and Skyrise only incur infrastructure costs. In this scenario, the recent Snowflake serverless offering is not cost-competitive with Athena.
@@ 4.3 Elasticity
[…]
We run the queries cold in 15 minute intervals to avoid any caching or provisioning effects in the AWS infrastructure. The results are shown in Figure 7. We report the aggregated runtimes of both queries. Given ideal elastic scalability, we would expect the query latency to be the same across all scale factors. We, however, observe that larger scale factors, from 10, but especially 1,000 and 10,000 are generally more prone to stragglers in function invocations and storage requests. The only configuration without straggler effects is scale factor 1, which runs on a single worker. Although Skyrise retriggers stragglers at different levels of the stack, detecting and mitigating them comes with an increasing overhead. This is the case specifically for the 10,000 configuration, which runs with a maximum of 2,500 query workers and suffers from IOPS limits in S3. Generally, without any upfront provisioning or load, Skyrise automatically scales the resources with the size of the query, and we see less than an order of magnitude (OOM) difference in query latency, despite the problem sizes varying by 5 OOMs. Thus, Skyrise effectively uses the elasticity inherent in its serverless compute and storage resources.
@@ 5.1 FaaS-based Data Processing
Following its initial public release in 2014, FaaS technology was notably adopted by Jonas et al. who developed PyWren [16]. PyWren is a prototypical implementation of a parallel programming model for serverless cloud environments. While PyWren targeted applications more generally, the potential of cloud functions for data analytics was shown by Kim et al. with Flint [53] and later Pu et al. with Locus [54]. Flint is a rewrite of Apache Spark’s execution layer using serverless building blocks like workers and disaggregated storage. In addition, Locus proposed a shuffling operator using S3 and AWS ElasticCache. Locus was extended by Zhang et al. with a serverless SQL query engine [55], called Caerus, focusing on analytical query processing. Skyrise is the first fully serverless system built entirely on serverless infrastructure without the need for servers hosting a driver, coordinator, or shuffle service. Having Skyrise open-source, allows for using it for research and teaching purposes at HPI and beyond.
The shuffling approaches from PyWren, Flint, and Corral [56] were enhanced by Lambada [22] and Starling [23] through staged shuffles, reducing the necessary storage requests for shuffles. Tiered shuffling was designed in Locus, Cackle [57], Pixels [58], Caerus /Jiffy [59] and Pocket [60]. Boxer [61] shuffles data via NAT and Flock [62] for streaming utilizes function invocation payloads [62]. Skyrise is the first system with tiered shuffling to hot serverless S3 Express storage.
Starling [23] addresses performance robustness when processing data on S3 through pipelining and solutions for I/O stragglers. Flint [63] employs a fine-granular fault-tolerance scheme with checkpointing and server selection to cope with cheap but transient servers. Salama et al. introduce a cost-based approach to selecting subsets of intermediate results to cache. In contrast, Skyrise tracks worker progress with query context and runtime statistics to prevent straggling workers due to computation or data skew.
@@
Combining stateful and stateless compute is an interesting new area that Pixels [58] and Cackle [57] introduced as hybrid compute to minize cost. For more cost-efficient serverless processing of high query volumes, Skyrise introduces partial result caching on serverless storage.